# A closer look at some cyclic semifields

### v1

@@ 3. Different generators of $\mathrm{Gal}(K/F)$ yield nonisomorphic Sandler algebras
we have $h(t)^{i}$ is well defined for all $i\leq m-1$, i.e., all $i$th powers of $h(t)$ are power-associative for $i\leq m-1$, and $h(t)h(t)^{m-1}=\sum_{i=0}^{m-1}\tau(a_{i})h(t)^{i}{\rm mod}_{r}g.$
@@ 4. Counting the non-isotopic nonassociative cyclic algebras of prime degree
(ii) Let $K/F$ be a Galois field extension of degree $n\geq 3$, and $\sigma_{1}$, $\sigma_{2}$ two generators of ${\rm Gal}(K/F)$. Suppose that $m\leq n$. Let $(Q,P,U)$ be an isotopy between two proper nonassociative algebras $K[t;\sigma_{1}]/K[t;\sigma_{1}](t^{m}-a_{1})$ and $K[t;\sigma_{2}]/K[t;\sigma_{2}](t^{m}-a_{2})$. Then there exists some $V\in{\rm Aut}(K)$ and some $h_{i}\in K^{\times}$ such that
$P=\left[\begin{array}[]{cccccc}VR(h_{1})&&...&\\ &VR(h_{2})&...&\\ &&VR(h_{3})&\\ ...&...&...&...\\ &&...&\\ &&...&VR(h_{m})\\ \end{array}\right].$
@@
for some $V\in{\rm Aut}(K)$ where $R(h_{ij})$ is right multiplication with $h_{ij}\in K$, as in [San1962, (34)]. Let $Q(1)=u_{0}+u_{1}t+\dots+t^{m-1}u_{m-1}$ for some $u_{i}\in K$ and $Q(t)=a_{0}+a_{1}t+\dots+t^{m-1}a_{m-1}$ for some $a_{i}\in K$. Analogously as in [San1962, (19)], we have $PL_{Q(z)}=L_{z}PL_{Q(1)}$. Put $z=t$ in $PL_{Q(z)}=L_{z}PL_{Q(1)}$, this way e obtain a set of linear equations identical to the ones in [San1962, (31)], just with the formula in the left hand side employing $\sigma_{1}$ and the ones on the right hand side $\sigma_{2}$. Arguing as in [San1962, p. 246], if exactly one element in the first row of $P$ is non-zero, then exactly one element in each row or column of $P$ is non-zero, and these will all lie on a transversal. We obtain hence analogously as in [San1962, p. 248] that
$P=\left[\begin{array}[]{cccccc}VR(h_{11})&&...&\\ &VR(h_{22})&...&\\ &&VR(h_{33})&\\ ...&...&...&...\\ &&...&\\ &&...&VR(h_{mm})\\ \end{array}\right]$
@@
$P=\left[\begin{array}[]{cccccc}VR(h_{1})&&...&\\ &VR(h_{2})&...&\\ &&VR(h_{3})&\\ ...&...&...&...\\ &&...&\\ &&...&VR(h_{m})\\ \end{array}\right].$
Without loss of generality we assume that $h_{1}=1$. This is possible, since for every $a\in K^{\times}$, $(id,D_{a},D_{a})$ is an autotopism of $A_{i}$, where $D_{a}$ denotes the diagonal matrix with $a$ as its diagonal entries. Hence we can multiply $P$ by a suitable autotopism to achieve $h_{1}=1$. As in [San1962, (44)], we also obtain that $Q=U=PL_{Q(1)}$. We continue along the lines of Sandler’s proof to see that also in our setting, $h_{k}=h_{1}\sigma_{2}(h_{1})\sigma_{2}^{2}(h_{1})\cdots\sigma_{2}^{k-1}(h_{1})$. Note that up to here, it is not required that the algebra is a division algebra. Then a short calculation analogous to the one given in [San1962, p. 264] shows that we obtain either that $\sigma(V(a_{1}))=a_{2}h_{1}\sigma_{2}(h)\sigma_{2}^{2}(h)\cdots\sigma_{2}^{m-1}(h)=a_{2}N_{K/F}(h)$ or $V(a_{1})=a_{2}h_{1}\sigma_{2}(h)\sigma_{2}^{2}(h)\cdots\sigma_{2}^{m-1}(h)=a_{2}N_{K/F}(h)$ for some $h\in K^{\times}$. In either case, this implies that $P$ is an isomorphism between the two algebras by Theorem 2.3. This means that $\sigma_{1}=\sigma_{2}$ by Theorem LABEL:sigmadistinct. ∎
@@
$\mathbb{F}_{q^{m}}[t;\sigma_{1}]/\mathbb{F}_{q^{m}}[t;\sigma_{1}](t^{m}-a_{1})$
@@
$\mathbb{F}_{q^{m}}[t;\sigma_{2}]/\mathbb{F}_{q^{m}}[t;\sigma_{2}](t^{m}-a_{2})$
@@
Corollary 4.5.
(i) Let $m$ not be prime and let $\varphi(m)$ be the number of generators of the cyclic Galois group ${\rm Gal}(K/F)$ of order $n$. Then we have $\varphi(m)$ different classes of non-isotopic semifields $(\mathbb{F}_{q^{m}}/\mathbb{F}_{q},\sigma_{i},a)$, one class for each choice of generator $\sigma_{i}$ of ${\rm Gal}(K/F)$.
(ii) If $m$ is prime and does not divide $q-1$, then there are exactly
$\frac{(q^{m}-q)(m-1)}{(q-1)m}$
non-isotopic Sandler semifields of order $q^{m^{2}}$ with center $F=\mathbb{F}_{q}$ and left, middle and right nucleus $K=\mathbb{F}_{q^{m}}$.
(iii) If $m$ is prime and divides $q-1$, then there are exactly
$(m-1)^{2}+(m-1)\frac{q^{m}-q-(q-1)(m-1)}{m(q-1)}$
non-isotopic Sandler semifields of order $q^{m^{2}}$ with center $F=\mathbb{F}_{q}$ and left, middle and right nucleus $K=\mathbb{F}_{q^{m}}$.
@@
We briefly discuss the $n=4$: Let $F$ be a finite field of odd prime characteristic. Let $K$ be the cyclic Galois field extension of degree four of $F$ with ${\rm Gal}(K/F)=\langle\sigma\rangle$. Then $K$ has exactly one quadratic subfield $E={\rm Fix}(\sigma^{2})$.
@@ 5. An explicit parametrization of nonassociative cyclic algebras in odd characteristic
Two proper semifields coordinatize the same non-Desarguesian plane if and only if they are isotopic. The subsequent results thus yield a full parametrization of the non-Desarguesian planes related to the semifields $(\mathbb{F}_{q^{m}}/\mathbb{F}_{q},\sigma,a)$.
@@ 5.1. When $m$ is an odd prime
Suppose $m$ is an odd prime and $\mathbb{F}_{q}$ contains a primitive $m$th root of unity $\zeta$. Then the distinct isomorphism classes of nonassociative cyclic algebras of degree $m$ over $F$ are represented by
@@
where:
  • $\mathbb{F}_{q^{m}}$ is the unique cyclic Galois field extensions of $\mathbb{F}_{q}$ of degree $m$;
@@ 6. What does this mean for the related linear codes?
Our results raise several questions that will be addressed in a future paper. It would certainly be interesting to see in how far our Second Main Theorem can be generalized to hold for other Petit algebras $K[t;\sigma]/K[t;\sigma]f$, when we still assume that the degree of $f$ is less or equal to the order $n$ of the automorphism $\sigma$ (the order may be infinite), but where $K/F$ does not need to be a Galois field extension. A straightforward generalization of the proof of Theorem 3.3 already shows:
@@
Let $F$ be an arbitrary field and let $n\geq 3$ be the degree of the cyclic Galois extension $K/F$. Let $\sigma_{1}$ and $\sigma_{2}$ be any two distinct generators of $\mathrm{Gal}(K/F)$, and suppose that $n\geq m$. Let
@@ 6.1. $(f,\sigma,\delta)$-codes
Skew polynomials $f\in R=K[t;\sigma,\delta]$ can be used to build linear codes (for a comprehensive exposition, cf. [BouLe2013]). Module $\sigma$-codes and $(f,\sigma,\delta)$-codes can be associated to the left ideals of a nonassociative Petit algebra $R/Rf$ for a suitable reducible $f\in R$ (choosing $\delta=0$ for the module $\sigma$-codes): the cyclic submodules studied in [BouUl2009, BUl2014, BUl2009] are exactly the left ideals in the Petit algebra $R/Rf$, the $(\sigma,\delta)$-codes of [BouLe2013] are the codes $\mathcal{C}$ associated to some left ideal of the Petit $R/Rf$. These ideals are generated by a right divisor $g$ of $f$.
@@
If $\mathcal{C}$ is a linear code over $\mathbb{F}_{q}$ of length $m$, then $\mathcal{C}$ is a $\sigma$-constacyclic code with constant $d$ if and only if the skew-polynomial representation $\mathcal{C}(t)$ with elements $a(t)$ obtained from $(a_{0},\dots,a_{m-1})\in\mathcal{C}$ is a left ideal of the Petit algebra $\mathbb{F}_{q}[t;\sigma]/\mathbb{F}_{q}[t;\sigma](t^{m}-d)$, which is generated by a monic right divisor $g$ of $f$ in $R$.
@@
Let $f(t)=t^{m}-d\in R=K[t;\sigma]$ and $F={\rm Fix}(\sigma)$. Then any right divisor $g$ of degree $k$ of $f=t^{m}-d$ can be used to construct a $\sigma$-constacyclic $[m,m-k]$-code with constant $d$.
Let $(K/F,\sigma,d)$ be a nonassociative cyclic division algebra of prime degree $m$, i.e. $f$ is irreducible. Therefore there are no $\sigma$-constacyclic codes with constant $d$, apart from the $[m,m]$-code associated with the nonassociative cyclic algebra $(K/F,\sigma,d)$ itself. So we need to look at algebras that are not division algebras in this context.
@@
Let $K=\mathbb{F}_{q^{n}}[t;\sigma]$ and $a_{i}\in K\smallsetminus F$. If $n\geq 3$ and $n>m$ then the Petit algebras $\mathbb{F}_{q}[t;\sigma]/\mathbb{F}_{q}[t;\sigma](t^{m}-a_{i})$ are nonassociative. Since
$K[t;\sigma_{1}]/K[t;\sigma_{1}](t^{m}-b)\not\cong K[t;\sigma_{2}]/K[t;\sigma_{2}](t^{m}-b)$

### v2

@@ 3. Different generators of $\mathrm{Gal}(K/F)$ yield nonisomorphic Sandler algebras
the $i$th power $h(t)^{i}$ is well defined for all $i\leq m-1$, i.e., all $i$th powers of $h(t)$ are power-associative for $i\leq m-1$, and $h(t)h(t)^{m-1}=\sum_{i=0}^{m-1}\tau(a_{i})h(t)^{i}{\rm mod}_{r}g.$
@@ 4. Counting the non-isotopic nonassociative cyclic algebras of prime degree
(ii) Let $K/F$ be a Galois field extension of degree $n\geq 3$, and $\sigma_{1}$, $\sigma_{2}$ two generators of ${\rm Gal}(K/F)$. Suppose that $n\geq m$. Let $(Q,P,U)$ be an isotopy between two proper nonassociative algebras $K[t;\sigma_{1}]/K[t;\sigma_{1}](t^{m}-a_{1})$ and $K[t;\sigma_{2}]/K[t;\sigma_{2}](t^{m}-a_{2})$. Then there exists some $V\in{\rm Aut}(K)$ and some $h_{i}\in K^{\times}$ such that
$P=\left[\begin{array}[]{cccccc}VR(h_{1})&0&...&0\\ 0&VR(h_{2})&0&...\\ 0&0&VR(h_{3})&...\\ ...&...&...&...\\ ...&...&...&0\\ 0&...&0&VR(h_{m})\\ \end{array}\right].$
@@
for some $V\in{\rm Aut}(K)$ where $R(h_{ij})$ is right multiplication with $h_{ij}\in K$, as in [San1962, (34)]. Let $Q(1)=u_{0}+u_{1}t+\dots+t^{m-1}u_{m-1}$ for some $u_{i}\in K$ and $Q(t)=a_{0}+a_{1}t+\dots+t^{m-1}a_{m-1}$ for some $a_{i}\in K$. Analogously as in [San1962, (19)], we have $PL_{Q(z)}=L_{z}PL_{Q(1)}$. Put $z=t$ in $PL_{Q(z)}=L_{z}PL_{Q(1)}$, this way we obtain a set of linear equations identical to the ones in [San1962, (31)], just with the formula in the left hand side employing $\sigma_{1}$ and the ones on the right hand side $\sigma_{2}$. Arguing as in [San1962, p. 246], if exactly one element in the first row of $P$ is non-zero, then exactly one element in each row or column of $P$ is non-zero, and these will all lie on a transversal. We obtain hence analogously as in [San1962, p. 248] that
$P=\left[\begin{array}[]{cccccc}VR(h_{11})&0&...&0\\ 0&VR(h_{22})&...&0\\ 0&0&VR(h_{33})&...\\ ...&...&...&...\\ ...&...&...&0\\ 0&...&0&VR(h_{mm})\\ \end{array}\right]$
@@
$P=\left[\begin{array}[]{cccccc}VR(h_{1})&0&...&0\\ 0&VR(h_{2})&0&...\\ 0&0&VR(h_{3})&...\\ ...&...&...&...\\ ...&...&...&0\\ 0&...&0&VR(h_{m})\\ \end{array}\right].$
Without loss of generality we assume that $h_{1}=1$. This is possible, since for every $a\in K^{\times}$, $(id,D_{a},D_{a})$ is an autotopism of $A_{i}$, where $D_{a}$ denotes the diagonal matrix with $a$ as its diagonal entries. Hence we can multiply $P$ by a suitable autotopism to achieve $h_{1}=1$. As in [San1962, (44)], we also obtain that $Q=U=PL_{Q(1)}$. We continue along the lines of Sandler’s proof to see that also in our setting, $h_{k}=h_{1}\sigma_{2}(h_{1})\sigma_{2}^{2}(h_{1})\cdots\sigma_{2}^{k-1}(h_{1})$. Note that up to here, it is not required that the algebra is a division algebra. Then a short calculation analogous to the one given in [San1962, p. 264] shows that we obtain either that $\sigma(V(a_{1}))=a_{2}h_{1}\sigma_{2}(h)\sigma_{2}^{2}(h)\cdots\sigma_{2}^{m-1}(h)=a_{2}N_{K/F}(h)$ or $V(a_{1})=a_{2}h_{1}\sigma_{2}(h)\sigma_{2}^{2}(h)\cdots\sigma_{2}^{m-1}(h)=a_{2}N_{K/F}(h)$ for some $h\in K^{\times}$. In either case, this implies that $P$ is an isomorphism between the two algebras by Theorem 2.3. This means that $\sigma_{1}=\sigma_{2}$ by [NevPu2024, Theorem 5.1]. ∎
@@
$\mathbb{F}_{q^{m}}[t;\sigma_{1}]/\mathbb{F}_{q^{m}}[t;\sigma_{1}](t^{m}-a_{1})=(\mathbb{F}_{q^{m}}/\mathbb{F}_{q},\sigma_{1},a_{1})$
@@
$\mathbb{F}_{q^{m}}[t;\sigma_{2}]/\mathbb{F}_{q^{m}}[t;\sigma_{2}](t^{m}-a_{2})=(\mathbb{F}_{q^{m}}/\mathbb{F}_{q},\sigma_{2},a_{2})$
@@
We briefly discuss the $n=4$ case: Let $F$ be a finite field of odd prime characteristic. Let $K$ be the cyclic Galois field extension of degree four of $F$ with ${\rm Gal}(K/F)=\langle\sigma\rangle$. Then $K$ has exactly one quadratic subfield $E={\rm Fix}(\sigma^{2})$.
@@ 5. An explicit parametrization of nonassociative cyclic algebras in odd characteristic
Two proper semifields coordinatize the same non-Desarguesian plane if and only if they are isotopic. The subsequent results thus yield a full parametrization of the non-Desarguesian planes related to the semifields $(\mathbb{F}_{q^{m}}/\mathbb{F}_{q},\sigma,a)$ of odd prime degree $m$.
@@ 5.1. When $m$ is an odd prime
Suppose $m$ is an odd prime, $\mathbb{F}_{q}$ has odd characteristic and contains a primitive $m$th root of unity $\zeta$. Then the distinct isomorphism classes of nonassociative cyclic algebras of degree $m$ over $F$ are represented by
@@
where $\mathbb{F}_{q^{m}}$ is the unique cyclic Galois field extensions of $\mathbb{F}_{q}$ of degree $m$;
  • 1.
@@ 6. What does this mean for the related linear codes?
Our results raise several questions that will be addressed in a future paper. It would certainly be interesting to see in how far our Second Main Theorem can be generalized to hold for other Petit algebras $K[t;\sigma]/K[t;\sigma]f$, when we still assume that the degree $m$ of $f$ satisfies that $n\geq m-1$ with $n$ the order of the automorphism $\sigma$ (the order may be infinite), but where $K/F$ does not need to be a Galois field extension. A straightforward generalization of the proof of Theorem 3.3 already shows:
@@
Let $F$ be an arbitrary field and let $n\geq 3$ be the degree of the cyclic Galois extension $K/F$. Let $\sigma_{1}$ and $\sigma_{2}$ be any two distinct generators of $\mathrm{Gal}(K/F)$, and suppose that $n\geq m-1$. Let
@@
6.1. Linear maximum rank distance codes
Linear maximum rank distance codes (MRD-codes) can be obtained from skew polynomial rings $K[t;\sigma,\delta]$, employing the invertible matrices representing the right multiplication in a Petit division algebra $K[t;\sigma,\delta]/K[t;\sigma,\delta]f$. Every proper nonassociative Petit division algebra hence canonically yields an $F$-linear MRD-code in $M_{mn}(K)$ of dimension $mn$, when $f$ has degree $m$. Since the left nucleus of a proper Petit division algebra is $K$, its right multiplication also canonically yields a set of invertible matrices in $M_{m}(K)$ that form an $F$-linear MRD-code.
Write $R(x)$ for the matrix of left multiplication by $x=x_{0}+x_{1}t+\cdots+x_{n-1}t^{n-1}\in(\mathbb{F}_{q^{m}}/\mathbb{F}_{q},\sigma,a)$, $x_{i}\in K$, with respect to the basis $\{1,t,\ldots,t^{n-1}\}$ in a division algebra $(\mathbb{F}_{q^{m}}/\mathbb{F}_{q},\sigma,a)$:
$R(x)=\left(\begin{array}[]{cccccc}x_{0}&a\sigma(x_{m-1})&...&a\sigma^{m-1}(x_{1})\\ x_{1}&\sigma(x_{0})&...&a\sigma^{m-1}(x_{2})\\ x_{2}&\sigma(x_{1})&...&a\sigma^{m-1}(x_{3})\\ ...&...&...&...\\ x_{m-2}&\sigma(x_{m-3})&...&a\sigma^{m-1}(x_{n-1})\\ x_{m-1}&\sigma(x_{m-2})&...&\sigma^{m-1}(x_{0})\\ \end{array}\right).$
These matrices define the MRD-code $\mathcal{C}_{\sigma,a}=\{R(x)\,|\,x\in A\}\subset M_{m}(\mathbb{F}_{q^{m}})$.
Theorem 6.2.
(i) The MRD-codes in $M_{m^{2}}(\mathbb{F}_{q})$ obtained from proper nonassociative Petit division algebras of the kind $\mathbb{F}_{q^{m}}[t,\sigma]/\mathbb{F}_{q^{m}}[t,\sigma](t^{m}-a)$ depend on the choice of the generator of the Galois group: the $\varphi(m)$ different possible choices of generators of ${\rm Gal}(\mathbb{F}_{q^{m}}/\mathbb{F}_{q})$ yield a total of $\varphi(m)$ different classes of non-equivalent $\mathbb{F}_{q}$-linear MRD-codes in $M_{m^{2}}(\mathbb{F}_{q})$ with minimum distance $m$.
(ii) The MRD-codes $\mathcal{C}_{\sigma,a}$ in $M_{m}(\mathbb{F}_{q^{m}})$ obtained from proper nonassociative Petit division algebras of the kind $\mathbb{F}_{q^{m}}[t,\sigma]/\mathbb{F}_{q^{m}}[t,\sigma](t^{m}-a)$ depend on the choice of the generator of the Galois group: different choices of generators of ${\rm Gal}(\mathbb{F}_{q^{m}}/\mathbb{F}_{q})$ yield a total of $\varphi(m)$ different classes of non-equivalent $\mathbb{F}_{q}$-linear MRD-codes $\mathcal{C}_{\sigma,a}$ in $M_{m}(\mathbb{F}_{q^{m}})$ with minimum distance $m^{2}$.
Proof.
(i) It is well-known that the isotopy classes of finite semifields of dimension $m^{2}$ over $F=\mathbb{F}_{q}$ are in one-one correspondence with the equivalence classes of $\mathbb{F}_{q}$-linear MRD-codes in $M_{m^{2}}(\mathbb{F}_{q})$ with minimum distance $m^{2}$. Our results show that different choices of generators of ${\rm Gal}(K/F)$ yield non-equivalent $\mathbb{F}_{q}$-linear MRD-codes with minimum distance $m^{2}$.
(ii) The isotopy classes of the proper nonassociative Petit algebras $\mathbb{F}_{q^{m}}[t,\sigma]/\mathbb{F}_{q^{m}}[t,\sigma](t^{m}-a)$ over $\mathbb{F}_{q}$ are in one-one correspondence with the equivalence classes of the $\mathbb{F}_{q}$-linear MRD-codes $\mathcal{C}_{\sigma,a}$ in $M_{m^{2}}(\mathbb{F}_{q})$ with minimum distance $m$, that are obtained employing the matrices that describe their right multiplication [She2019]. Thus again different choices of generators yield non-equivalent $\mathbb{F}_{q}$-linear MRD-codes with minimum distance $m$. ∎
Since these codes are special cases of the codes constructed in [She2019] we conjecture that at least some of the other equivalence classes of Sheekey’s codes also will depend on the choice of generator of the Galois group involved.
Theorem 6.3.
(i) Suppose $m$ is an odd prime, $\mathbb{F}_{q}$ has odd characteristic and contains a primitive $m$th root of unity $\zeta$. Then the distinct equivalence classes of $\mathbb{F}_{q}$-linear MRD-codes in $M_{m^{2}}(\mathbb{F}_{q})$ with minimum distance $m^{2}$, respectively, the distinct equivalence classes of $\mathbb{F}_{q}$-linear MRD-codes $\mathcal{C}_{\sigma,a}$ in $M_{m}(\mathbb{F}_{q^{m}})$ with minimum distance $m$, that correspond to the algebras $(\mathbb{F}_{q^{m}}/\mathbb{F}_{q},\sigma,a)$ can be obtained by using the (unique) cyclic Galois field extension $\mathbb{F}_{q^{m}}$ of $\mathbb{F}_{q}$ of degree $m$, choosing a generator $\sigma$ of $\mathrm{Gal}(\mathbb{F}_{q^{m}}/\mathbb{F}_{q})$, and $a\in\mathcal{S}(\mathbb{F}_{q^{m}})$.
(ii) Suppose $\mathbb{F}_{q^{2}}=\mathbb{F}_{q}(\sqrt{c})$ for some $c\in\mathbb{F}_{q}^{\times}\smallsetminus(\mathbb{F}_{q}^{\times})^{2}$ and $\mathrm{Gal}(\mathbb{F}_{q^{2}}/\mathbb{F}_{q})=\langle\sigma\rangle$. Then the distinct equivalence classes of $\mathbb{F}_{q}$-linear MRD-codes in $M_{4}(\mathbb{F}_{q})$ with minimum distance $4$, respectively, the distinct equivalence classes of $\mathbb{F}_{q}$-linear MRD-codes $\mathcal{C}_{\sigma,a}$ in $M_{2}(\mathbb{F}_{q^{m}})$ with minimum distance $2$, that that correspond to the algebras $(\mathbb{F}_{q^{2}}/\mathbb{F}_{q},\sigma,a)$, can be obtained by choosing $a$ from the set
or, alternatively, from
Proof.
(i) By Theorem 5.1 and [San1962, Theorem 9], we have a parametrization of the non-isotopic semifields of order $q^{m^{2}}$ obtained via the construction $(\mathbb{F}_{q^{m}}/\mathbb{F}_{q},\sigma,a)$ when ${\rm Fix}(\sigma)=F=\mathbb{F}_{q}$, with one representative for each isotopy class. Isotopy classes of finite semifields of dimension $m$ over $F=\mathbb{F}_{q}$ are in one-one correspondence with the equivalence classes of $\mathbb{F}_{q}$-linear MRD-codes in $M_{m^{2}}(\mathbb{F}_{q})$ with minimum distance $m^{2}$, respectively, with the equivalence classes of $\mathbb{F}_{q}$-linear MRD-codes in $M_{m}(\mathbb{F}_{q^{m}})$ with minimum distance $m$. Thus we have one representative for each corresponding MRD-code.
(ii) is proved as (i) but employing Theorem 5.2. ∎
@@ 6.2. $(f,\sigma,\delta)$-codes
Skew polynomials $f\in R=K[t;\sigma,\delta]$ can be used to build linear codes (for a comprehensive “historical” exposition, cf. [BouLe2013]). Module $\sigma$-codes, $(f,\sigma,\delta)$-codes and the more recent skew polycyclic codes [BP2025] all can be associated to the left ideals of a suitable nonassociative Petit algebra $R/Rf$ for a suitable reducible $f\in R$: the cyclic submodules studied in [BouUl2009, BUl2014, BUl2009] are exactly the left ideals in the Petit algebra $R/Rf$, the $(\sigma,\delta)$-codes of [BouLe2013] are the codes $\mathcal{C}$ associated to some left ideal of the Petit algebra $R/Rf$. These ideals are generated by a right divisor $g$ of $f$.
@@
If $\mathcal{C}$ is a linear code over $\mathbb{F}_{q^{n}}$ of length $m$, then $\mathcal{C}$ is a skew $\sigma$-constacyclic code with constant $a$ if and only if the skew-polynomial representation $\mathcal{C}(t)$ with elements $a(t)$ obtained from $(a_{0},\dots,a_{m-1})\in\mathcal{C}$ is a left ideal of the nonassociative Petit algebra $\mathbb{F}_{q^{n}}[t;\sigma]/\mathbb{F}_{q^{n}}[t;\sigma](t^{m}-a)$, which is generated by a monic right divisor $g$ of $f$ in $R$.
@@
Let $f(t)=t^{m}-a\in K[t;\sigma]$ and $F={\rm Fix}(\sigma)$. Then any right divisor $g$ of degree $k$ of $f=t^{m}-a$ can be used to construct a skew $\sigma$-constacyclic $[m,m-k]$-code with constant $a$. So we need to look for Petit algebras $K[t,\sigma]/K[t,\sigma](t^{m}-a)$ that are not division algebras in this context.
@@
Let $R=\mathbb{F}_{q^{n}}[t;\sigma]$ and $a_{i}\in\mathbb{F}_{q^{n}}\smallsetminus\mathbb{F}_{q}$. The Petit algebras $\mathbb{F}_{q^{n}}[t;\sigma]/\mathbb{F}_{q^{n}}[t;\sigma](t^{m}-a_{i})$ are proper nonassociative algebras. Assume that $n\geq 3$ and $n\geq m-1$. Since
$\mathbb{F}_{q^{n}}[t;\sigma_{1}]/\mathbb{F}_{q^{n}}[t;\sigma_{1}](t^{m}-b)\not\cong\mathbb{F}_{q^{n}}[t;\sigma_{2}]/\mathbb{F}_{q^{n}}[t;\sigma_{2}](t^{m}-b)$